OpenStudy (anonymous):

Medal to whoever answers it!!!! Solve for x: -5(x - 3) = 30 x = 9 x = -9 x = 3 x = -3

OpenStudy (anonymous):

-3

OpenStudy (texaschic101):

- 5(x - 3) = 30 -- distribute the -5 through the parenthesis first -5x + 15 = 30 -- subtract 15 from both sides -5x + 15 - 15 = 30 - 15 -- combine like terms -5x = 15 -- divide by -5 to get x by itself -5x/-5 = 15/-5 x = - 3 Do you have any questions ? I tried to do it step by step so you could understand how it is done